# A Joint Learning Framework for Bridging Defect Prediction and Interpretation

### v1

@@ Abstract
Over the past fifty years, numerous software defect prediction (SDP) approaches have been proposed. However, the ability to explain why predictors make certain predictions remains limited. Explainable SDP has emerged as a promising solution by using explainable artificial intelligence (XAI) methods to clarify the decision-making processes of predictors. Despite this progress, there is still significant potential to enhance the reliability of existing approaches. To address this limitation, we treat defect prediction and the corresponding interpretation as two distinct but closely related tasks and propose a joint learning framework that allows for the simultaneous training of the predictor and its interpreter. The novelty of our approach lies in two main aspects: 1. We design feedback loops that convey the decision-making logic from the predictor to the interpreter. This ensures a high level of conciseness in decision logic and feature engineering for both the predictor and the interpreter, enabling the interpreter to achieve reliable local and global interpretability. 2. We incorporate the interpretation results as a penalty term in the loss function of the joint-learning framework. This not only improves the accuracy of the predictor but also imposes a stronger constraint on the reliability of the interpreter. We validated our proposed method against several existing explainable SDPs across multiple datasets. The results demonstrate its effectiveness in both interpretation and defect prediction. The source code for the proposed method is available at: https://github.com/BugPredictor/software-defect-prediction.git
@@ I Introduction
For nearly half a century, software defect prediction (SDP) has been a significant research focus in the field of software engineering [1, 2]. Its primary aim is to identify defective code early in the software development process, helping practitioners prioritize quality assurance tasks, especially when resources are limited or deadlines are tight [49]. Inspired by the success of machine learning (ML) in various disciplines, such as computer vision and machine translation, researchers have explored applying different ML algorithms to improve the accuracy of SDP. However, despite numerous efforts, existing SDP models often fail to provide actionable guidance for practitioners, making practical application challenging [51].
Recent studies have highlighted that the limited applicability of SDP in real-world scenarios is closely linked to the lack of interpretability [3, 4]. Interpretability refers to the ability to provide comprehensible insights into the decision-making logic for humans. Without it, developers struggle to understand the logic behind SDP predictions, leading to distrust and reluctance to adopt these techniques [3, 5]. Explaining prediction results can be particularly challenging for ML-based predictors, especially those using deep learning (DL). The complex structures and large number of parameters in such models obscure straightforward interpretation.
To address this issue, researchers have investigated the feasibility of explainable SDP, incorporating explainable artificial intelligence (XAI) into SDP. Efforts in explainable SDP are generally categorized into two groups [6]: global and local interpretations. Global approaches focus on explaining the overall behavior of a defect predictor, offering insights into how the model makes decisions across all instances [7]. Local approaches provide clear evidence for why a given instance is predicted as buggy or clean. The prediction results provide actionable guidance for practitioners to address key questions, such as why a specific prediction was made and how to mitigate associated risks [57]. As a result, practitioners often prefer to use local interpretation techniques in practice [3]. Nevertheless, local techniques often encounter challenges with reliability under different settings [52].
@@
Figure 1: Interpretations of LIME for a code fragment.
To illustrate the challenges associated with local approaches, we take the interpretation for ”AbstractCommand.java” from the Apache ActiveMQ project as an example. The explanations generated by LIME with different samplers (i.e., SMOTE and RUS) are shown in Fig. 1 LABEL:sub@fig_1:LIME_RUS. The explanations generated by applying LIME twice to the same file are shown in Fig. 1 LABEL:sub@fig1:LIME_two. These figures present the rankings of features that contribute to the predictions. Features displayed in red indicate their contribution to predicting buggy code, while those displayed in green indicate their contribution to the prediction of clean code.
In Fig. 1 LABEL:sub@fig_1:LIME_RUS, we can see that although the SDP model with different samplers predicts the file as buggy, their explanations vary significantly. Among the ten features selected by LIME for the SMOTE-based predictor, only two are also selected for the RUS-based predictor. Fig. 1 LABEL:sub@fig1:LIME_two further demonstrates that the two explanations generated by running the same predictor twice show considerable variation. For instance, in the first explanation, seven features contribute to the prediction of buggy code, whereas in the second explanation, the number increases to eight.
To address this issue, we consider defect prediction and the corresponding interpretation as two distinct but strongly related tasks and propose a joint-learning framework in this paper. Specifically, we design multiple feedback loops between the interpreter and predictor to simultaneously achieve reliable interpretation and high prediction accuracy. The specific contributions are as follows:
  • To improve the reliability of interpretation, we revised the loss function of the knowledge distillation (KD) framework and introduced output fidelity loss and feature fidelity loss. This strategy constructs a knowledge transfer loop between the interpreter and predictor, ensuring a high level of conciseness in decision-making logic and feature engineering power.
  • To improve the accuracy of the predictor, we incorporate the interpretation results as a penalty term within the loss function of the joint-learning framework. This enables the predictor to capture more discriminative features guided by the interpretation results while imposing stronger restrictions on the fidelity of the interpreter to the predictor.
  • Extensive experiments are conducted on 20 datasets with an in-depth analysis. The results indicate that the joint-learning framework improves both the reliability of interpretation and the accuracy of the predictor.
The paper is organized as follows: Section II provides an overview of the existing research on SDP and interpretation methods used in SDP. In Section III, we describe our proposed method in detail, while in Section IV, we provide the details of the experiment. In Section V,we discuss the threats to validity. Finally, in Section VI, we present our conclusion and future works.
@@ II Related Work
This section provides an overview of SDP and explores existing techniques for explainable SDP.
@@ II-A Software defect prediction
The primary goal of SDP is to classify code fragments (e.g., method or class) as either clean or buggy. Constructing a typical SDP model involves three steps. The first step is to transform software metrics, such as cohesion, coupling, complexity, encapsulation, inheritance, and size, into corresponding numerical representations [10]. The second step labels code fragments as buggy or clean based on post-release defects. These defects can be collected from a Bug Tracking System (BTS) by connecting bug reports to their bug-fixing activities. Code fragments associated with these bug-fixing activities are regarded as buggy, while all others are labeled as clean. Finally, the labeled code fragments serve as training data for an ML-based classifier, which is then used to predict whether code fragments in a later version are buggy or clean.
@@
Numerous ML algorithms, including Fuzzy self-organizing maps, K-means, SVM, random forest, logistic regression, etc., have been extensively employed in previous research [15, 16, 17, 18, 19, 20, 21]. However, metrics used in SDP consistently exhibit correlations and redundancies [11], compromising the effectiveness of traditional ML-based predictors [12]. To address this issue, various DL models have been investigated to improve SDP performance. The initial DL-based defect predictor, Deeper, introduced by Yang et al. [22], employed a deep belief network to identify discriminative information in the input data. Experimental results showed that Deeper could identify 32.22% more defects than many traditional ML-based approaches. Inspired by Yang’s work, many researchers have proposed different DL-based solutions. Specifically, Hoang et al. proposed a convolutional neural network (CNN)-based defect prediction model [23], which experimental results indicated outperformed Deeper. Qiao et al. proposed another DL-based model called DPNN to predict the number of defects [24]. Experimental results showed a 14% reduction in mean square error and an 8% increase in the squared correlation coefficient. Li et al. proposed another CNN-based defect predictor, DP-CNN [25], which achieved better performance than other methods. As demonstrated, DL has significantly improved the performance of SDP.
@@
To ensure the generalization, we employ the DP-CNN proposed in [25] as the predictor and evaluate the difference in accuracy before and after integrating it into the joint-learning framework.
II-B Interpretation of defect prediction
In many practical application scenarios, practitioners often ask: What is the underlying decision logic of the predictor? What is the reason for predicting a code fragment as buggy? How do different metrics influence the prediction? These questions highlight an urgent requirement for explainable defect prediction. However, highly accurate SDP models (e.g., Random Forest, DL) are often black-box in nature, making them difficult to interpret.
@@ II-B Interpretation of defect prediction
Efforts to improve the interpretability of predictions can be broadly divided into two main categories [6]: global and local approaches. The global approaches offer a high-level view of how inputs influence the model’s predictions and reveal patterns that are consistent across the model’s entire operation, rather than focusing on individual predictions. Many ML algorithms (e.g., decision trees, logistic regression techniques) and statistical methods (e.g., ANOVA, variable importance) fall into this category[26, 27]. These methods share two common aspects: 1. Simulatability: Every step of the decision-making process is comprehensible. For example, decision trees are simulatable since we can observe the state of every node as well as the feature thresholds used to separate samples. 2. Transparency: The decision-making process can be broken down into interpretable steps. For instance, the interpretability of a decision tree is defined by the computational path from the root node to the leaf nodes. However, global approaches cannot provide detailed interpretations for each code fragment, and their relatively simple structure often fails to guarantee optimal predictive accuracy. Consequently, many practical applications have adopted local methods.
@@
In contrast to global approaches, local methods such as LIME [8] and BreakDown [28], usually interpret the prediction of each code fragment as a list of ordered features. Feedback from practitioners indicates that the interpretations produced by local methods can offer more actionable guidance [52]. However, recent studies have reported that local techniques lack reliability under various conditions [30, 31, 52]. Firstly, the interpretations of the same code can be inconsistent when different data sampling techniques are used. In SDP, sampling is frequently used due to the scarcity of buggy codes. It is an effective way to mitigate the effects of imbalanced data on the ML algorithms [53]. In [52], Jiho Shin et al., investigated the consistency of local methods across five commonly used data sampling techniques. Their analysis revealed that both the features and the ranking of features in the interpretation vary significantly when different data sampling techniques are employed. Secondly, the interpretations of a code fragment do not exhibit consistency when multiple interpretations are executed. The underlying intuition of local methods regards the interpreter $g$ and the predictor $f$ as two independent ML algorithms and leverages $g$ to approximate the behavior of $f$ within the local area of an instance $x$, denoted as $\pi_{x}$. In general, local methods cannot guarantee that the $\pi_{x}$ required for multiple interpretations of the same instance are the same. So different $\pi_{x}$ can lead to distinct interpretations for the same code fragment [31]. Finally, the predictions produced by the interpreter and the predictor for the same code fragment are not consistent. To reduce the difficulty of interpretation, many local methods define the interpreter as a simple-structure ML model. Taking the most widely used local interpretation technique LIME as an example, it sets a linear regression model as the interpreter, and the interpretation of SDP is oversimplified as a linear transformation between input metrics and prediction outcomes. Recently, [31] emphasized that attempting to comprehend an intricate model by employing a simple model might be overly optimistic. The inconsistency between the predictions generated by the interpreter and the predictor indicates that the interpretation fails to capture the underlying decision-making logic of the predictor.
@@ II-C Knowledge distillation
Knowledge distillation (KD) is a typical model compression technique that transfers knowledge from a teacher model (e.g., DL models) to a student model (e.g., a shallow neural network) [32, 33]. KD is regarded as a promising framework for improving the reliability of interpretation due to the faithful knowledge transfer channel between the student and teacher models. In the health care domain, [34] leveraged distilled knowledge to discover disease patterns. Similarly, in the business domain, [35] distilled dynamical pricing knowledge from a complex black-box DL model, optimizing revenue while maintaining interpretability. In this paper, we designate the predictor as the teacher model and the interpreter as the student model, and transfer the decision-making knowledge of predictor to interpreter based on the KD principle. The main difference between our approach and existing methods lies in the revised loss functions of KD framework, which enable us to train the interpreter and predictor in a collaborative manner.
@@ III Proposed Method
Fig. 2 provides an overview of our approach, which consists of four modules: metric selection, predictor and interpreter design, joint learning framework, and local interpretation.
@@ III-A Metric selection
We carefully selected 16 metrics for our method after a thorough investigation of existing literature and metric suites, including the works of [10, 36, 37]. The details of metrics are presented in Table I. The selection of these metrics was guided by two primary criterias, ensuring their relevance and effectiveness in the context of defect prediction: (1) Expressiveness: The metrics needed to be sufficiently expressive, enabling them to capture subtle variations in code quality, structure, and complexity between similar pieces of source code. (2) Proven Effectiveness: Each selected metric has been validated in prior studies [17, 22, 38, 39], demonstrating its utility in software defect prediction tasks. By adhering to these criterias, we constructed a dataset that covered a broad spectrum of code attributes, ranging from complexity measures to process-related characteristics.
@@
CC McCabe’s cyclomatic complexity
@@ III-B1 Defect predictor
To ensure the generality of our method, we chose a DL-based model, DP-CNN [25] as the predictor. It consists of three main components: the convolution layer, max pooling layer, and classification layer. To further improve accuracy, we added an attention layer to the original model. The overall architecture of DP-CNN is presented in Fig. 3.
@@
The working procedure of DP-CNN can be summarized as follows: First, the input defect metrics are fed into the convolution layer to extract correlations between features. Then feature maps are then passed through the max pooling layer, which reduces the dimensionality while retaining the most significant features. To further enhance the model’s ability to focus on critical patterns, the attention layer dynamically assigns weights to the feature maps. Finally, the processed features are passed through a fully connected neural network and a softmax classifier to generate the prediction result. The correlations between metrics are crucial for defect prediction. Two characteristics of the convolution layer, sparse connectivity and shared weights, make it particularly suitable for this task [25]. Sparse connectivity enables CNNs to have a wider receptive field, capturing non-linear correlations between metrics. Shared weights ensure that each CNN kernel has the same parameters, allowing predictors to capture these correlations regardless of their position in the input metrics. The max pooling layer retains the most significant features by selecting the maximum local element in the convolutional output, enhancing the model’s robustness to changes in the position of metrics in the input. The attention layer dynamically assigns weights to the inputs, providing a flexible way to capture correlations between feature maps. The fully connected neural network is responsible for converting the feature maps generated by previous layers into a logit. The softmax block is positioned at the end of the entire neural network, transforming the logits into the probabilities.
@@ III-B2 Interpretation model
Following the KD principle, we employ the soft decision tree (SDT) as the interpreter [40]. The SDT is a variant of the fuzzy decision tree, sharing a similar structure with the ordinary decision tree but differing in the definition of nodes. In SDT, nodes are defined as neurons with learnable weight $W$ and bias $b$. Weights reflect the importance of each feature in the decision-making process at the node. The output of each node is defined as $P_{i}(x_{i})=\sigma(x_{i}W_{i}+b_{i})$, which determines the probability of transitioning to the right subtree. The $\sigma(\cdot)$ denotes the activation function and $x_{i}$ is the input of the $i$-th node. The output at the $l$-th leaf is defined as:
@@
The reasons we use SDT as an interpreter are as follows: Owing to its neural-like node architecture, SDT exhibits comparable feature engineering capabilities to CNN, enabling it to accurately retain the decision-making knowledge of CNN. Furthermore, in comparison to CNN, the complexity of SDT has been reduced, rendering the decision-making process more comprehensible.

### v2

@@ Abstract
Understanding why defect predictors classify software components as defective or clean is essential for software engineers that helps identify the root causes of defects and develop actionable bug-fixing plans. Existing solutions employ various explainable artificial intelligence (XAI) methods to clarify the decision-making processes of defect predictors. However, these post-hoc explanation techniques have two main limitations: 1. the interpretation results do not accurately reflect the model’s decision logic, and 2. they do not contribute to improving the performance of defect predictors. To address these limitations, we treat defect prediction and its corresponding interpretation as two distinct but closely related tasks, proposing a joint learning framework that trains the predictor and its interpreter simultaneously. The novelty of our approach lies in two main aspects: 1. we design a feedback loop that transfers decision logic from the predictor to the interpreter, ensuring a high degree of conciseness for both components. 2. we incorporate interpretation results as a penalty term in the loss function of the joint learning framework. This not only enhances the accuracy of the predictor but also proposes a stronger constraint on the reliability of the interpreter. We validate our method against several existing explainable software defect predictors across multiple datasets. The results demonstrate its effectiveness. The source code of our method is available at: https://github.com/BugPredictor/software-defect-prediction.git
@@ I Introduction
Software defect prediction (SDP) has attracted significant attention in software engineering for over five decades [1, 2]. The primary goal of this research is to identify defective code early in the software development process, helping practitioners prioritize quality assurance tasks, especially when resources are limited or deadlines are tight [3].
In recent years, inspired by the success of machine learning (ML) and deep learning (DL) in various disciplines, researchers have introduced different ML/DL algorithms into SDP. Despite significant improvements in the performance of SDP models, many practitioners remain hesitant to integrate SDP techniques into practical applications [4]. Such a situation is primarily due to the lack of interpretability of these models. In practical scenarios, practitioners often ask: What is the reason for predicting a code fragment as buggy? How do different metrics influence the prediction? These questions underscore an urgent requirement for explainability in SDP [5, 6, 7].
To address the interpretation issue, researchers have proposed various explainable SDP models that integrate different explainable artificial intelligence (XAI) techniques with SDP. Existing explainable SDP models are generally categorized into two groups [8]: global and local approaches. Global approaches focus on explaining the overall behavior of a defect predictor, offering insights into how the model makes decisions across all instances [9]. In contrast, local approaches provide clear evidence for why a given instance is predicted as buggy or clean. Compared to global approaches, local approaches focus on identifying the root cause of specific defects and provide actionable guidance for practitioners to address key questions, such as why a specific prediction was made and how to mitigate associated risks [10]. As a result, local interpretation techniques are generally considered more favorable in practice [5, 11].
However, the task properties of interpretation and defect prediction are quite different, making explainable SDP a challenging problem. On one hand, SDP relies on high-complexity ML/DL models to capture intricate relationships among software metrics and defect patterns. These models are often considered as black boxes, with decision-making processes that are difficult to interpret. On the other hand, XAI methods favor simpler models that provide a clear, human-understandable logical reasoning path. Bridging these two competing objectives—complexity and interpretability—remains a significant challenge. Most existing solutions adopt a post-hoc explanation strategy, in which an SDP model is first trained and subsequently interpreted by XAI techniques [12, 13]. Although this strategy offers a degree of interpretability, it presents two main challenges: 1. the interpretations often fail to reliably reflect the model’s actual decision logic, and 2. they do not contribute to improving the predictive performance of defect models.
@@
Figure 1: An example of unreliable interpretations.
To better illustrate the challenges associated with local approaches, we examine the interpretation of ”AbstractCommand.java” from the Apache ActiveMQ project as an example. The explanations generated by LIME using different data sampling techniques (i.e., SMOTE and RUS) are shown in Fig. 1 LABEL:sub@fig_1:LIME_RUS, and the explanations generated by executing LIME twice are shown in Fig. 1 LABEL:sub@fig1:LIME_two. The features highlighted in red contribute to the prediction of buggy code, while those highlighted in green contribute to the prediction of clean code. From the Fig. 1 LABEL:sub@fig_1:LIME_RUS, we observe that the explanations vary significantly. Among the ten features selected by SMOTE-based LIME, only two are also selected by RUS-based LIME. Fig. 1 LABEL:sub@fig1:LIME_two further illustrates the issue of poor reliability. In the first round of explanation, seven features are identified as correlated with buggy code, whereas in the second round, this number increases to eight.
To address this issue, we treat defect prediction and the corresponding interpretation as two distinct yet closely related tasks, proposing a joint learning framework in this paper. Specifically, we design multiple feedback loops between the interpreter and predictor to simultaneously achieve reliable interpretation and high prediction accuracy. The main contributions are as follows:
  • To enhance the reliability of interpretation, we introduce two novel loss functions: output fidelity loss and feature fidelity loss. This strategy establishes a knowledge transfer loop between the interpreter and predictor, ensuring both a high level of conciseness in decision-making logic and feature engineering power between these two components.
  • To improve the accuracy of the predictor, we incorporate the interpretation results as a penalty term in the loss function of the joint learning framework. This strategy enables the predictor to capture more discriminative features guided by the interpretation results while simultaneously imposing stronger constraints on the fidelity of the interpreter to the predictor.
  • Extensive experiments are conducted on twenty datasets with an in-depth analysis. The results demonstrate that the joint learning framework improves both the reliability of interpretation and the accuracy of the predictor.
The remainder of paper is organized as follows: Section II provides an overview of the existing research on SDP and interpretation methods used in SDP. Section III describes our proposed method in detail, while Section IV presents the experimental setup and results. Section V discusses the threats to validity. Finally, Section VI presents our conclusions and outlines directions for future work.
@@ II-A Software defect prediction
Constructing an SDP model involves three steps. The first step is to transform software metrics, such as cohesion, coupling, complexity, into numerical representations [14]. The second step labels code fragments as buggy or clean based on post-release defects. Finally, leverage the labeled code fragments to train an ML/DL-based classifier for identifying future defect-prone code.
@@
Over the past fifty years, numerous ML algorithms, including fuzzy self-organizing maps, K-means, SVM, Random Forest, and logistic regression, have been extensively employed in SDP research [15, 16, 17, 18, 19, 20, 21]. However, software metrics often exhibit correlations and redundancies [22], and traditional ML algorithms struggle to capture the complex dependencies among them [23]. To address this issue, various DL models have been introduced in SDP. One of the earliest explorations, Deeper [24], employed a deep belief network to identify discriminative information in input data. Experimental results showed that Deeper identified 32.22% more defects than many traditional ML-based approaches. Inspired by the Deeper, researchers have introduced a sophisticated DL algorithms into SDP to enhance performance and broaden application scenarios. For example, [25, 26] leaverage convolutional neural networks (CNNs) to build defect prediction models, achieving better performance than other DL-based approaches. Additionally, Qiao et al. employed a DL model to predict the number of defects [27]. Experimental results showed a 14% reduction in mean squared error and an 8% increase in squared correlation coefficient.
@@
To ensure generalizability, we employ DP-CNN, proposed in [26], as the predictor and evaluate the accuracy differences before and after integrating it into the joint learning framework.
II-B Explainable defect prediction
@@
Efforts in explainable SDP can be broadly divided into two main categories [8]: global and local approaches. The global approaches provide a high-level view of how inputs features influence the model’s predictions. Many ML algorithms (e.g., decision trees, logistic regression) and statistical methods (e.g., ANOVA, variable importance) fall into this category[28, 29]. However, global approaches cannot offer detailed interpretations for individual code fragment, and their simple structure often fails to ensure optimal predictive accuracy. As a result, many practical applications have adopted local methods instead.
@@
Local interpretation methods such as LIME [12] and BreakDown [30], primarily focus on interpreting specific code fragments. However, recent studies have reported that local techniques often lack reliability under various conditions [31, 32, 11]. Firstly, interpretations can be inconsistent when different data sampling techniques are used or when the same interpreter is executed multiple times. Local interpretation methods typically employ an interpreter to approximate the predictor’s behavior within a local area of the given instance. The local area is generated by the sampling algorithm. However, the randomness of the sampling process makes it difficult to maintain consistency across local areas—regardless of whether the same or different sampling methods are used [11]. As a result, the same code fragment may yield distinct interpretations[32]. Secondly, interpretations tend to be oversimplified. Local interpretation methods usually employ ML algorithms with simple structure as interpreters. For example, LIME utilizes a linear regression model as the interpreter. The interpretation of SDP is oversimplified as a linear transformation between input metrics and prediction outcomes. Recent study [32] reported that attempting to comprehend an intricate model by employing a simple model might be overly optimistic. Such interpretations often fail to capture the underlying decision-making logic of the predictor.
@@ II-C Knowledge distillation
Knowledge distillation (KD) is a model compression technique that transfers knowledge from a teacher model (e.g., DL models) to a student model (e.g., a shallow neural network) [33, 34]. It is considered a viable solution for enhancing the reliability of interpretation due to the knowledge transfer mechanism between the teacher and student models. For instance, [35] utilized distilled knowledge to identify disease patterns, while [36] distilled dynamic pricing knowledge from a complex black-box DL model. In this paper, we treat the predictor and the interpreter as the teacher and student models, respectively, and design knowledge transfer channels based on the KD principle. The key distinction of our approach lies in the collaborative training mechanism: unlike existing KD methods, which keep the parameters of the teacher model fixed, our method trains both the interpreter and the predictor in a collaborative manner.
@@ III Proposed Method
Fig. 2 provides an overview of our proposed approach, which consists of four modules: metric selection, predictor and interpreter design, joint learning framework, and local interpretation.
@@ III-A Metric selection
We selected sixteen metrics for our method based on a review of existing literature [14, 37, 38]. The details of these metrics are presented in Table I. The selection was guided by two criterias: (1) expressiveness: The metrics needed to be sufficiently expressive to capture subtle variations in code quality, structure, and complexity. (2) effectiveness: Each selected metric has been demonstrated its utility in SDP tasks [17, 24, 39, 40]. According to these criteria, we constructed a dataset that encompasses a broad spectrum of code attributes, ranging from complexity measures to process-related characteristics.
@@
$C$ McCabe’s cyclomatic complexity
@@ III-B1 Defect predictor
In this paper, we chose a DL-based model, DP-CNN [26] as the predictor. It consists of three main components: a convolution layer, a max pooling layer, and a classification layer. To further improve predictive accuracy, we extend the original model by integrating an attention mechanism. The overall architecture of DP-CNN is illustrated in Fig. 3.
@@
The DP-CNN model predicts defect-proneness through the following process: Selected metrics are first fed into the convolutional layer to capture correlations among features. The output is then passed through the max pooling layer to highlight key features while reducing redundancy. Next, the attention layer focus on critical defect patterns by dynamically assigning weights. Finally, a fully connected network followed by softmax classifier generates the prediction. We choose the CNN-based model as the predictor for three key reasons: (1) Sparse connectivity enables a wider receptive field to capture non-linear correlations. (2) Weight sharing ensures positional invariance in feature extraction [26]. (3) Max pooling enhances robustness by preserving significant features, while the attention layer further refines feature importance.
@@ III-B2 Interpretation model
Following the KD principle, we employ the soft decision tree (SDT) as the interpreter [41]. The SDT is a variant of the fuzzy decision tree, sharing a similar structure with the ordinary decision tree but differing in the definition of nodes. In SDT, nodes are defined as neurons with learnable weights $W$ and biases $b$. The output of each node is defined as $P_{i}(x_{i})=\sigma(x_{i}W_{i}+b_{i})$, which determines the probability of transitioning to the right subtree. The $\sigma(\cdot)$ denotes the activation function and $x_{i}$ is the input of the $i$-th node. The output at the $l$-th leaf node is defined as:
@@
The reasons we use SDT as an interpreter are as follows: Owing to its neural-like node structure, SDT exhibits comparable feature engineering capabilities to CNN, enabling it to simulate the decision-making knowledge of CNN. Furthermore, the complexity of SDT has been reduced, making the decision-making process more comprehensible.